# MergeMix: A Unified Augmentation Paradigm for Visual and Multi-Modal Understanding

### v1

@@ Abstract
Vision–language alignment in multi-modal large language models (MLLMs) typically relies on supervised fine-tuning (SFT) or reinforcement learning (RL). SFT is stable and efficient but requires large-scale human annotations and cannot capture subtle preferences, while RL brings in a reward signal for training, but suffers from overhead and instability. These limitations highlight a trade-off between scalability, robustness, and alignment quality. To address this, we propose MergeMix, a training-time augmentation paradigm that bridges SFT and RL. It first applies an attention-aware image mixing via token merge with more cluster representation and spatial context, and then presents a preference-driven training paradigm for MLLMs by building preference pairs with mixed images and raw images, and optimizing via SimPO loss. As a mixup augmentation, MergeMix enhances attention consistency and efficiency, surpassing other heuristic-based methods in classification. Extensive experiments demonstrate that MergeMix achieves competitive accuracy with improved efficiency, providing a scalable approach to preference alignment in classification and MLLMs.
@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) (Liu et al., 2024b; Bai et al., 2025; Tong et al., 2024) have recently demonstrated remarkable capabilities in integrating visual and textual information, enabling a wide range of applications from visual question answering to multi-modal reasoning. Since MLLMs are typically pre-trained on massive web-scale datasets, forcing them to possess a wide range of knowledge and general reasoning capabilities, Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL)-based preference optimization (Yang et al., 2025b) have emerged as two primary paradigms for aligning MLLMs with human preferences and specific task requirements. However, SFT depends on high-quality instruction–response annotations and optimizes the likelihood of reference responses, which does not explicitly model relative preferences between outputs. RL-based methods such as RLHF are more preference-aware, but they require an additional reward model that may encode bias or be exploited by reward.
@@
Due to the shortcomings of data quality and training efficiency, some works (Zhu et al., 2024; 2025; Luo et al., 2024; Tan et al., 2025; Wang et al., 2024) try to build performance pairs for optimization. How to build the preference pair with control and high-quality data for model training is the remaining open question. For example, SeVa (Zhu et al., 2024) proposed a preference optimization method by building a loser through some classic augmentation (i.e., RandomCrop). Then, select the different responses for optimizing the model by a DPO loss (Rafailov et al., 2023). However, these methods have two drawbacks: the augmentations are highly random, and the DPO loss cannot be related to the data, which means SeVa can only select useful training data. Those technical causes SeVa can not control the quality of the loser, which is harmful for some visual question answering tasks, and reduces the training data by selecting “hard negatives”. Hence, we investigate an interesting question: Is it necessary to propose novel techniques rather than some classical machine learning methods in the MLLM scenario?
@@
In this paper, we revisit the mixup augmentation, which synthesizes mixed samples and corresponding labels with given mixing ratios. However, two main challenges arise as illustrated in Figure 1: (1) achieving an optimal trade-off between efficiency and performance of mixup augmentations that rely on saliency-based metrics, (2) extending the augmentation to MLLMs properly, from classical image corruptions to data-dependent samples. Motivated by these perspectives, we propose a novel training framework called MergeMix, which builds preference pairs for MLLM training through data augmentation methods and ranking loss, thereby bridging the gap between SFT and RL. Figure 2 shows the two scenarios of MergeMix. (a) We introduce MergeMix, a novel data augmentation that generates mixed samples through token merge techniques. A bipartite soft matching gathers the similarity information that brings the context, making the mask retain useful features. Meanwhile, MergeMix links the merge ratio and mixing ratio, aligning the information density of samples, enabling precise mixing data generation. (b) We propose a preference-driven SFT paradigm for MLLMs, where augmented samples are regarded as non-preferred responses (Loser) and clean samples as preferred responses (Winner). This enables preference optimization via SimPO loss (Meng et al., 2024) without relying on reward models. Altogether, Figure 1 shows these contributions yield an efficient and effective training strategy that achieves stronger alignment with human preferences while preserving the stability and scalability of SFT. Since the optimization object has a direct relationship with augmentation, it obtains a more robust ability in calibration. Extensive experiments show that MergeMix, as a training-time augmentation paradigm, achieves competitive performance in both image classification and MLLM benchmark with favorable efficiency.
@@
    We use the naive token merge to obtain a local clustered attention map and recover the initial shape by source map, which can generate the mixed image with more continuous features.
@@
    We design a training paradigm for supervised fine-tuning of MLLMs by generating augmented images as losers, building preference pairs with raw images, and optimization via ranking loss.
@@
Figure 1: Efficiency and performance for MergeMix: (a) The training time vs. accuracy of mixup methods with the DeiT-Small model. (b) The image classification Top-1 accuracy vs. training epochs of different mixup methods on the CIFAR100 dataset with the DeiT-Tiny model. (c) The radar plot of the results on part VQA tasks by LLaVA-7B, LLaVA with SFT, and MergeMix.
@@ 2 Related Work
The Mixup method mitigates model overfitting by generating augmented samples through mixing two different images within a mini-batch. Broadly, Mixup methods can be categorized into two types: Static, which relies on human priors or randomness, and Adaptive, a data-dependent type that leverages certain metrics to guide the mixing process. (i) Static: MixUp (Zhang et al., 2017) generates mixed samples via linear interpolation with $\lambda$. CutMix (Yun et al., 2019) extends this idea from the global pixel level to a local patch level by constructing a mask of size proportional to $\lambda$ to mix images. ResizeMix (Qin et al., 2020) ensures that features from at least one class are always preserved in the mixed sample by resizing the source image before mixing. Other methods, e.g., FMix (Harris et al., 2020), SmoothMix (Lee et al., 2020), GridMix (Baek et al., 2021), and StarMix (Jin et al., 2024b), focus on improving the mask to obtain more suitable mixed samples. (ii) Adaptive: SaliencyMix (Uddin et al., 2021) employs a saliency extractor to identify informative patches in images for mixing. Attentive-CutMix (Walawalkar et al., 2020) and SuperMix (Dabouei et al., 2021) utilize a teacher model to guide mask generation. PuzzleMix (Kim et al., 2020) and Co-Mix (Kim et al., 2021) generate appropriate masks based on gradient information obtained from a forward pass of the samples. AutoMix (Liu et al., 2022) and AdAutoMix (Qin et al., 2024a) adopt an end-to-end bi-optimization paradigm to produce mixed samples. TransMix (Chen et al., 2022), SMMix (Chen et al., 2023), and MixPro (Zhao et al., 2023) specifically enhance ViTs by computing attention scores from a forward pass to generate feature-aware masks and further refine the label ratio through attention scores.
@@
Token Merging (Bolya et al., 2023) proposes to merge similar tokens by Key similarity for ViT-based models to achieve efficiency and acceleration. In MLLMs, images and texts will incur a significant number of tokens, which are often full of redundant information. Obvious researchers bring the token compression into MLLMs. Overall, we divide the methods that reduce tokens into 2 types, Reduce in Encoder and Reduce in Decoder. (i) Reduce in Encoder: MADTP (Cao et al., 2024) aims to achieve MLLM acceleration by purging visual tokens. LLaVA-PruMerge (Shang et al., 2024) uses the attention of [CLS] token to select clustering centers and then merges the remaining tokens with lower attention through a KNN clustering and weighted clustering center updating mechanism. VisionZip (Yang et al., 2025a), instead, retains visual tokens with high attention scores and subsequently merges the remaining tokens through clustering. Others, such as TokenPacker (Li et al., 2025), AVG-LLaVA (Lan et al., 2025), MustDrop (Liu et al., 2024c), and LLaVolta (Chen et al., 2024a), achieve acceleration by choosing a metric to sample TopK visual tokens. FastVLM (Vasu et al., 2025) proposes an Efficient Vision Encoder to reduce visual tokens. (ii) Reduce in Deocder: PyramidDrop (Xing et al., 2024) divides the token compression process in LLM into multiple stages, which employs a pyramidal token drop to avoid losing too much visual information in shallower layers. ATP-LLaVA (Ye et al., 2025) proposes an Adaptive Token Pruning (ATP) module that reduces the number of tokens in the decoder layer. ZipVL (He et al., 2024) proposes a dynamic ratio allocation strategy via the importance token, adaptively determined based on the distribution of attention scores in a particular layer, rather than a fixed hyperparameter.
@@ 3 Preliminaries
We define $\mathbb{X}$ to be the set of training samples and $\mathbb{Y}$ the set of ground truth of the corresponding labels. For each sample pair $(x,y)$, we randomly sample two pairs in $\mathbb{X}$ and $\mathbb{Y}$, with $\lambda$ in $\text{Beta}(\alpha,\alpha)$, the mixed images and labels are generated by applying the optimized mask $\mathcal{M}$ and ratio $\hat{\lambda}$, which come from a defined policy $\mathcal{P}(\cdot,\cdot,\cdot)$ according to Eq. (1):
@@
where the $\odot$ denotes element-wise multiplication. Policy $\mathcal{P}(\cdot,\cdot,\cdot)$ aims for the $\mathcal{M}$ to retain more features in the mixed sample. The $\hat{\lambda}$ keeps the initial sampling mixing ratio when the policy is without optimization; otherwise, the $\lambda$ can be re-computed by some metrics.
@@
where $(x,y^{+})$ and $(x,y^{-})$ denote the preferred and dispreferred responses respectively, and $s_{\theta}$ is a scoring function that reflects model preference. Different approaches (e.g., PPO (Schulman et al., 2017), DPO (Rafailov et al., 2023)) instantiate $\pi_{\theta}$ in various ways, but all share the same principle.
@@
where $\pi_{\theta}$ denotes the policy model (in our case, an MLLM such as LLaVA-v1.5 (Liu et al., 2024b) or Qwen-VL (Bai et al., 2025)), and $\pi_{\text{ref}}$ represents a frozen reference model used to preserve alignment with the original pre-trained distribution. $\sigma$ is the sigmoid function, and $\beta>0$ is a temperature-like scaling factor that controls the sharpness of preference separation. Intuitively, DPO encourages the policy to assign higher likelihood to preferred responses ($y^{+}$) than to non-preferred ones ($y^{-}$), while maintaining proximity to the reference model.
@@ 4.1 Image Mixing via Token Merge
In MergeMix, we leverage the relationship between the merge ratio and mixing ratio. The merge ratio measures the information of raw samples, while the mixing ratio balances the information between mixing samples, thereby enabling precise data generation of mixed inputs and labels. In this subsection, we first introduce MergeMix on the input space. Then we use the designed mixing policy $\mathcal{P}(,\cdot,)$ to obtain the mixed images $\hat{x}$ with the mask $\mathcal{M}$.
@@
Since token merge aggregates non-similar tokens into compact representations $Z_{K}$, the resulting matrix preserves local feature structures more effectively. In contrast, the vanilla TopK selection adopts a greedy sampling strategy with linear complexity $\mathcal{O}(N)$, which discards low-ranked tokens directly and thus loses spatial relationships. Alternatively, the Bipartite Soft Matching (BSM) approach performs global pairwise matching with quadratic complexity, yielding a more balanced and globally optimal merging of tokens. To reconstruct the full-resolution attention map, we introduce a recovering function $\mathcal{R}_{K\to L}(\cdot,\cdot)$ that expands the merged attention map $A_{K}$ back to its original length $A_{L}$ according to Eq. (6):
@@ 4.2 A Unified Augmentation Paradigm: From Image Classification to MLLMs
Figure 3: Overall illustration of MergeMix for MLLM. MergeMix performs attention-based mask mixing guided by the ToMe Vision Encoder, recovering token attention scores and generating a mixed image through an augmenter. Specifically, Token Merging hierarchically merges visual tokens through Bipartite Soft Matching (BSM) and merges to enhance efficiency. The model is trained with both the SFT and ranking losses.
Under this optimization objective, the role of the mixing ratio $\lambda$ is to serve as a metric that quantifies the presence of feature information from the two samples. While this metric cannot directly reflect the true characteristics of the data, certain adaptive methods can constrain the model to generate mixed samples where the mixing ratio progressively approximates the target value (Jin et al., 2024a). Some works, like LUMix (Sun et al., 2022), DecoupleMix (Liu et al., 2023), and SUMix (Qin et al., 2024b), use a defined policy for some hand-crafted mixup methods, and find that it is more efficient than optimizing a better mask way.
Since we introduce a token merge technology that inherently enables information aggregation and selection, the entire training process of the model requires consideration of not only simple spatial ratios but also the degree of information integration within the model. So, we proposed a Gaussian-based sampling to refine the ratio, where the merged tokens and the mask values jointly control the mean and std. This smooth transition directly alleviates changes from linear mapping and yields more robust augmentations, with its formulation given in Eq. (8):
@@
where the $\mathcal{N}(\cdot,\cdot)$ denotes a Gaussian function, $\mu$ and $\sigma$ represent the merged ratio and mixing ratio repetitively. $\tau$ as a hyperparameter, set to 1e-5. Then, we obtain the re-scaled mixing ratio $\hat{\lambda}$ with spatial and ToMe model inherent features to optimize the model when training. In total, the loss of mixup training as Eq. (9):
@@
LLaVA (Liu et al., 2024b) uses a standard conditional language modeling loss for SFT. In MLLMs, we are given an instruction–response pair $(x,y)$, where $x$ denotes the multi-modal data, inducing vision and text, and $y=(y_{1},y_{|y|})$ denotes the target response. The SFT loss is defined as:
@@
This objective needs to maximize the likelihood of GT responses, aiming to align the data.
In Section 3, we introduced the DPO loss, which can be decomposed into two components: the SFT part and the ranking optimization part. In our approach, we replace the ranking component with SimPO Meng et al. (2024), where $y$ denotes the target sequence (response) and $|y|$ denotes its length. Furthermore, Since $\lambda$ reflects information similarity between augmented and raw image (interpreted as “loser degree” in MLLMs), we link it to $\gamma\to 1-\lambda$: larger $\lambda$ represent higher similarity and harder discrimination, reduces $\gamma$ to avoid over-optimization on trivial differences; smaller $\lambda$ represent greater dissimilarity and easier tasks increases $\gamma$ to strengthen constraints for clearer preference distinction. The mixed SimPO loss replacement is Eq. (11):
$\mathcal{L}^{\text{Mix}}_{\mathrm{SimPO}}=-\mathbb{E}_{(x,\hat{x},y)\sim\mathcal{D}}\Big[\log\sigma\big(\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)-\frac{\beta}{|y|}\log\pi_{\theta}(y\mid\hat{x})-(1-\lambda)\big)\Big].$ (11)
This reformulated loss strictness with sample difficulty, enabling more robust preference optimization. Finally, the total loss of our training paradigm is Eq. (12):

### v2

@@ Abstract
Vision-language alignment in multi-modal large language models (MLLMs) relies on supervised fine-tuning (SFT) or reinforcement learning (RL). To align multi-modal large language models (MLLMs) in the post-training stage, supervised fine-tuning (SFT) is a stable choice but requires human annotations and lacks task generalizations, while Reinforcement Learning (RL) searches for better answers from reward signals but suffers from computational overhead and instability. To achieve balance among scalability, efficiency, and alignment generalizations, we propose MergeMix, a unified paradigm that bridges SFT and RL with an efficient Token Merge based Mixup augmentation. As for the Mixup policy, we generate contextual aligned mixed images with the corresponding labels according to the merged attention maps with cluster regions. Then, we enhance the preference-driven paradigm for MLLMs by building preference pairs with raw images and MergeMix-generated ones and optimizing the soft preference margin with the mixed SimPO loss. Extensive experiments demonstrate that MergeMix not only achieves dominant classification accuracy as an augmentation method but also improves generalization abilities and alignment of MLLMs, providing a new learning paradigm for preference alignment with training efficiency and stability.
@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) (liu2024llava; bai2025qwen2.5; tong2024cambrian) have recently demonstrated remarkable capabilities in integrating visual and textual information, enabling a wide range of applications from visual question answering to multi-modal reasoning. Since MLLMs are typically pre-trained on massive web-scale datasets, forcing them to possess a wide range of knowledge and general reasoning capabilities, Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL)-based preference optimization (yang2025visionthink) have emerged as two primary paradigms for aligning MLLMs with human preferences and specific task requirements. However, SFT depends on high-quality instruction–response annotations and optimizes the likelihood of reference responses, which does not explicitly model relative preferences between outputs. RL-based methods such as RLHF are more preference-aware, but require an additional reward model that may introduce bias or be exploited by the reward signal.
@@
Due to the shortcomings of data quality and training efficiency, some works (zhu2024seva; zhu2025nsft; luo2024probing; tan2025beyond; wang2024sima) try to build performance pairs for optimization. How to build the preference pair with control and high-quality data for model training is the remaining open question. For example, SeVa (zhu2024seva) proposed a preference optimization method by building a loser through some classic augmentation (i.e., RandomCrop). Then, select the different responses for optimizing the model by a DPO loss (rafailov2023dpo). However, these methods have two drawbacks: the augmentations are highly random, and the DPO loss cannot be related to the data, which means SeVa can only select useful training data. Those technical causes SeVa can not control the quality of the loser, which is harmful for some visual question answering tasks, and reduces the training data by selecting “hard negatives”. Hence, we investigate an interesting question: Is it necessary to propose novel techniques rather than some classical machine learning methods in the MLLM scenario?
@@
In this paper, we revisit the mixup augmentation, which synthesizes mixed samples and corresponding labels with given mixing ratios. However, two main challenges arise as illustrated in Figure 1: (1) achieving an optimal trade-off between efficiency and performance of mixup augmentations that rely on saliency-based metrics, (2) extending the augmentation to MLLMs properly, from classical image corruptions to data-dependent samples. Motivated by these perspectives, we propose a novel training framework called MergeMix, which builds preference pairs for MLLM training through data augmentation methods and ranking loss, thereby bridging the gap between SFT and RL. Figure 2 shows the two scenarios of MergeMix. (a) We introduce MergeMix, a novel data augmentation that generates mixed samples through token merge techniques. A bipartite soft matching strategy captures similarity information that preserves contextual features, ensuring the mask retains useful information. Meanwhile, MergeMix links the merge ratio and mixing ratio, aligning mixed images with the corresponding labels, enabling precise mixing data generation. (b) We propose a preference-driven paradigm for MLLMs, where augmented samples are defined as non-preferred responses (Loser) and clean samples as preferred responses (Winner). This paradigm facilitates preference tuning via the mixed SimPO loss, and leverages the mixing ratio as the soft preference margin to enable adaptive optimization. Altogether, Figure 1 shows these contributions yield an efficient and effective training strategy that achieves stronger alignment with human preferences while preserving the stability and scalability of SFT. Since the optimization object has a direct relationship with augmentation, it obtains a more robust ability in calibration. Extensive experiments show that MergeMix, as a training-time augmentation paradigm, achieves competitive performance in both image classification and MLLM benchmark with favorable efficiency.
@@
    We use token merging to obtain a local clustered attention map, enabling the generation of mixed images with cluster regions, a label re-scaling strategy aligned the mixed images with their corresponding labels, achieve well performance on both overhead and classification accuracy.
@@
    We enhance the preference tuning paradigm for supervised fine-tuning of MLLMs, where mixed images are treated as losers, the mixing ratio is used as a soft preference reward score, and optimize the model adaptively via the mixed SimPO loss.
@@
Figure 1: Efficiency and for MergeMix: (a) The training time vs. accuracy of mixup methods with the DeiT-Small model. (b) The image classification Top-1 accuracy vs. training epochs of different mixup methods on the CIFAR100 dataset with the DeiT-Tiny model. (c) The radar plot of the results on part VQA tasks by LLaVA-7B, LLaVA with SFT, and MergeMix.
@@ 2 Related Work
The Mixup method mitigates model overfitting by generating augmented samples through mixing two different images within a mini-batch. Broadly, Mixup methods can be categorized into two types: Static, which relies on human priors or randomness, and Adaptive, a data-dependent type that leverages certain metrics to guide the mixing process. (i) Static: MixUp (zhang2017mixup) generates mixed samples via linear interpolation with $\lambda$. CutMix (yun2019cutmix) extends this idea from the global pixel level to a local patch level by constructing a mask of size proportional to $\lambda$ to mix images. ResizeMix (qin2020resizemix) ensures that features from at least one class are always preserved in the mixed sample by resizing the source image before mixing. Other methods, e.g., FMix (harris2020fmix), SmoothMix (lee2020smoothmix), GridMix (Baek2021gridmix), and StarMix (jin2024starlknet), focus on improving the mask to obtain more suitable mixed samples. (ii) Adaptive: SaliencyMix (uddinsaliencymix) employs a saliency extractor to identify informative patches in images for mixing. Attentive-CutMix (icassp2020attentive) and SuperMix (dabouei2021supermix) utilize a teacher model to guide mask generation. PuzzleMix (kim2020puzzle) and Co-Mix (kim2021co) generate appropriate masks based on gradient information obtained from a forward pass of the samples. AutoMix (liu2022automix) and AdAutoMix (qin2024adversarial) adopt an end-to-end bi-optimization paradigm to produce mixed samples. TransMix (chen2022transmix), SMMix (chen2023smmix), and MixPro (zhao2023mixpro) specifically enhance ViTs by computing attention scores from a forward pass to generate feature-aware masks and further refine the label ratio through attention scores. DiffuseMix (islam2024diffusemix) and GenMix (islam2024genmix) generate mixed samples by a diffusion model (zhu2025obs) for label preserving.
@@
Token compression methods (bolya2023tome; tao2025dycoke; shao2025holitom; chen2025streamingtom; tao2025omnizip) mainly propose to merge or drop redundant tokens to achieve efficiency and acceleration. In MLLMs, images and texts will incur a significant number of tokens, which are often full of redundant information. Obvious researchers bring the token compression into MLLMs. Overall, we divide the methods that reduce tokens into 2 types, Reduce in Encoder and Reduce in Decoder. (i) Reduce in Encoder: MADTP (cao2024madtp) aims to achieve MLLM acceleration by purging visual tokens. LLaVA-PruMerge (shang2024llavapruerge) uses the attention of [CLS] token to select clustering centers and then merges the remaining tokens with lower attention through a KNN clustering and weighted clustering center updating mechanism. VisionZip (yang2025visionzip), instead, retains visual tokens with high attention scores and subsequently merges the remaining tokens through clustering. Others, such as TokenPacker (li2025tokenpacker), AVG-LLaVA (lan2025avg), MustDrop (liu2024mustdrop), and LLaVolta (chen2024LLaVolta), achieve acceleration by choosing a metric to sample TopK visual tokens. FastVLM (vasu2025fastvlm) proposes an Efficient Vision Encoder to reduce visual tokens. (ii) Reduce in Deocder: PyramidDrop (xing2024pyramiddrop) divides the token compression process in LLM into multiple stages, which employs a pyramidal token drop to avoid losing too much visual information in shallower layers. ATP-LLaVA (ye2025atp) proposes an Adaptive Token Pruning (ATP) module that reduces the number of tokens in the decoder layer. ZipVL (he2024zipvl) proposes a dynamic ratio allocation strategy via the importance token, adaptively determined based on the distribution of attention scores in a particular layer, rather than a fixed hyperparameter.
@@ 3 Preliminaries
We define $\mathbb{X}$ to be the set of training samples and $\mathbb{Y}$ the set of ground truth of the corresponding labels. For each sample pair $(x,y)$, we randomly sample two pairs in $\mathbb{X}$ and $\mathbb{Y}$, with $\lambda$ in $\text{Beta}(\alpha,\alpha)$. The mixed images and labels are generated by applying the optimized mask $\mathcal{M}$ and ratio $\hat{\lambda}$, which come from a defined policy $\mathcal{P}(\cdot,\cdot,\cdot)$ according to Eq. (1):
@@
where the $\odot$ denotes element-wise multiplication. Policy $\mathcal{P}(\cdot,\cdot,\cdot)$ aims for the $\mathcal{M}$ to retain more features in the mixed sample. The $\hat{\lambda}$ keeps the initial sampling ratio when without optimization; otherwise, the $\lambda$ can be re-computed by some metrics (i.e., MergeMix uses the total mask values).
@@
where $(x,y^{+})$ and $(x,y^{-})$ denote the preferred and dispreferred responses respectively, and $s_{\theta}$ is a scoring function that reflects model preference. Different approaches (e.g., PPO (schulman2017ppo), DPO (rafailov2023dpo)) instantiate $\pi_{\theta}$ in various ways, but all share the same principle.
@@
where $\pi_{\theta}$ denotes the policy model (in our case, an MLLM such as LLaVA-v1.5 (liu2024llava) or Qwen-VL (bai2025qwen2.5)), and $\pi_{\text{ref}}$ represents a frozen reference model used to preserve alignment with the original pre-trained distribution. $\sigma$ is the sigmoid function, and $\beta>0$ is a temperature-like scaling factor that controls the sharpness of preference separation. Intuitively, DPO encourages the policy to assign a higher likelihood to preferred responses ($y^{+}$) than to non-preferred ones ($y^{-}$), while maintaining proximity to the reference model.
@@ 4.1 Image Mixing via Token Merge
In MergeMix, we leverage the relationship between the merge ratio and mixing ratio. The merge ratio measures the information of raw samples, while the mixing ratio balances the information between mixing samples, thereby enabling precise data generation of mixed inputs and labels. In this subsection, we first introduce MergeMix on the input space. Then we use the designed mixing policy $\mathcal{P}(\cdot,\cdot,\cdot)$ to obtain the mixed images $\hat{x}$ with the mask $\mathcal{M}$.
@@
Since token merging aggregates non-similar tokens into compact representations $Z_{K}$, the resulting matrix preserves local feature structures more effectively. In contrast, the vanilla TopK selection adopts a greedy sampling strategy with linear complexity $\mathcal{O}(N)$, which discards low-ranked tokens directly and thus loses spatial relationships. Alternatively, the Bipartite Soft Matching (BSM) approach performs global pairwise matching with quadratic complexity, yielding a more balanced and globally optimal merging of tokens. To reconstruct the full-resolution attention map, we introduce a recovering function $\mathcal{R}_{K\to L}(\cdot,\cdot)$ that expands the merged attention map $A_{K}$ back to its original length $A_{L}$, which shows in Fig 5 and according to Eq. (6):
@@
Figure 3: Overall illustration of MergeMix for MLLM. (a) MergeMix performs attention-based mask mixing guided by the ToMe Vision Encoder, recovering token attention scores and generating a mixed image through an augmenter. Specifically, Token Merging hierarchically merges visual tokens via Bipartite Soft Matching (BSM) to enhance efficiency, which is trained with both the SFT and ranking losses. (b) Case study of preference data generated by MergeMix with LLaVA-v1.5-7B.
@@ 4.2 A Unified Augmentation Paradigm: From Image Classification to MLLMs
Under this optimization objective, the role of the mixing ratio $\lambda$ is to serve as a metric that quantifies the presence of feature information from the two samples. While this metric cannot directly reflect the true characteristics of the data, certain adaptive methods can constrain the model to generate mixed samples where the mixing ratio progressively approximates the target value (jin2024survey). Some works, like LUMix (sun2022lumix), DecoupleMix (liu2023decoupledmix), and SUMix (qin2024sumix), use a defined policy for some hand-crafted mixup methods, and find that it is more efficient than optimizing a better mask way.
[…]
@@
where the $\mathcal{N}(\cdot,\cdot)$ denotes a Gaussian function, $\mu$ and $\sigma$ represent the merged ratio and mixing ratio repetitively. $\tau$ as a hyperparameter, set to 1e-5. Then, we obtain the re-scaled mixing ratio $\hat{\lambda}$ with spatial and ToMe model inherent features to optimize the model when training. In total, the loss of mixup training as:
@@
LLaVA (liu2024llava) uses a standard conditional language modeling loss for SFT. In MLLMs, we are given an instruction–response pair $(x,y)$, where $x$ denotes the multi-modal data, inducing vision and text, and $y=(y_{1},y_{|y|})$ denotes the target response. The SFT loss is defined as:
@@
This objective needs to maximize the likelihood of GT responses, aiming to align the data. In Section 3, we introduced the DPO loss, which can be decomposed into two components: the SFT part and the ranking optimization part. In our approach, we replace the ranking component with SimPO (meng2024simpo), where $y$ denotes the target sequence (response) and $|y|$ denotes its length. Furthermore, since $\lambda$ reflects information similarity between augmented and raw image (interpreted as “loser degree” in MLLMs), we link it to $\gamma\to 1-\hat{\lambda}$: larger $\lambda$ represent higher similarity and harder discrimination, reduces $\gamma$ to avoid over-optimization on trivial differences; smaller $\lambda$ represent greater dissimilarity and easier tasks increases $\gamma$ to strengthen constraints for clearer preference distinction. The mixed SimPO loss replacement is Eq. (11):
$)].\displaystyle\mathcal{L}^{\text{Mix}}_{\mathrm{SimPO}}=-\mathbb{E}_{(x,\hat{x},y)\sim\mathcal{D}}\Big[\log\sigma\big(\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)-\frac{\beta}{|y|}\log\pi_{\theta}(y\mid\hat{x})-(1-\hat{\lambda}$$\big)\Big$]. (11)
This reformulated loss strictness with sample difficulty, enabling more robust preference optimization. Finally, the total loss of our training paradigm is written as: